### v6

@@ III-B Dual-activated lightweight attention ResNet50
A new approach for automatic breast cancer classification is proposed in this research using a new version of the ResNet50 technique called Dual-activated lightweight attention ResNet50 (DALAResNet50). The DALAResNet50 approach incorporates ResNet50, Global pooling, channel recalibration, Dual Activation functions, and the newly designed fully connected layer into a unified approach to overcome the drawbacks of the traditional ResNet50 model. The pre-trained ResNet50 is first utilized as a baseline architecture. Subsequently, the attention module is integrated into the fourth residual block of ResNet50 to pay attention to the small-scale nuclei cells of breast cancer. The attention mechanism improves the ability of feature learning by extracting the silent semantic information to overcome the class imbalance issue. The proposed attention module is characterized by the amalgamation of adaptive global pooling and a two-layer fully connected network, employing LeakyReLU [10] and ReLU [11] as dual activation functions. The traditional ResNet50 uses conventional pooling layers to reduce the spatial dimensions of the feature maps while retaining important information. However, the conventional pooling layers operate over local regions of the input feature maps, reducing their spatial dimensions but discarding spatial information, making the ResNet50 sensitive to spatial translations of the input feature map and increasing the number of parameters in the network. Therefore, this research proposed average-based global pooling that calculates the average value of each feature map across all spatial locations, effectively summarizing each feature’s presence throughout the entire feature map. The main advantage of global pooling is that it provides a fixed-size representation of the whole feature map, making the ResNet50 less sensitive to spatial translations of the input and reducing the total number of parameters in the network. Additionally, the DALAResNet50 adopts the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero. This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks. The ReLU activation function [11] is further chosen for the second fully connected network to improve the model performance and image feature representation. Combining these two activation functions enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for breast cancer image classification tasks. Finally, the newly added fully connected layer is positioned at the end of the DALAResNet50 approach to address the complexity of binary classification problems.
@@
Fig. 2 and Fig. 3 illustrate the flowcharts of the standard ResNet approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. A comprehensive comparative analysis of these differences is detailed in Table 2. The visualization of the DALAResNet50 approach and the entire technical framework are depicted in Fig. 4 and Fig. 5.
@@ IV-A BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed dataset statistics in Table 3 cover sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 6 presents these pathological images as a practical example.
Refer to caption 40x
Refer to caption 100x
Refer to caption 200x
Refer to caption 400x
@@
Benign
625 644 623 588
Malignant
1370 1437 1390 1232
Total No of Images
1995 2081 2013 1820
@@ IV-C Experiment implementation
In deep learning, the configuration of various parameters is paramount. Notably, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention. In addition, setting the training batch size and the number of iterations is significant. In addition, setting the training batch size and the number of iterations is substantial. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to balance training effectiveness and computational efficiency. Finally, this study also specifically considered the dropout rate setting. Previous studies generally set the dropout rate between 0.25 and 0.5. In processing the unbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability. Therefore, this study chose a lower dropout rate of 0.25.
The method proposed in this study, and the benchmark model illustrated in the Table apply the above settings to ensure the experiments’ fairness and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
@@
SEResNet50 [39]
0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
DensNet121 [40]
0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
VGG16 [41]
0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
VGG16Inception [42]
0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
ViT [43]
0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
Swin-Transformer [44]
0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
Dinov2_Vitb14 [45]
0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
RseNet50 [9]
0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
@@
SEResNet50 [39]
0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
DensNet121 [40]
0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
VGG16 [41]
0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
VGG16Inception [42]
0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
ViT [43]
0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
Swin-Transformer [44]
0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
Dinov2_Vitb14 [45]
0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
RseNet50 [9]
0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
@@
SEResNet50 [39]
2696.22 2814.30 2711.89 2356.18 27073090 2644.65
DenseNet121 [40]
2352.46 2643.79 2423.36 2078.34 9028970 2347.49
VGG16 [41]
3212.95 3182.54 3011.53 2640.19 134268738 3011.80
VGG16Inception [42]
3074.99 3253.50 3303.12 2682.61 38139938 3078.56
ViT [43]
2932.11 3022.54 3013.86 2691.16 51375106 2914.92
Swin-Transformer [44]
3020.11 3132.42 2940.98 2592.96 86745274 2921.62
Dinov2_Vitb14 [45]
2895.02 2303.25 2245.98 2017.15 86582018 2365.35
ResNet50 [9]
2920.53 2933.26 2307.94 2097.33 21797672 2564.77
@@ IV-D Baseline selection
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze unbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another.”
@@ IV-E Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically. The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly [52]. IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance. This approach offers a more holistic assessment of the model’s performance, encompassing sample diversity and fairness [53]. GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
Table 6 showcases the performance of the model on the test is set across various image resolutions, such as 40X, 100X, 200X, and 400X. Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean. These results demonstrate that the network model has a significant advantage regardless of which image resolution is used. This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
Table 7 compares the model with eight other network models regarding the number of model parameters and convergence speed. Notably, the DenseNet-121 model possesses the fewest parameters. However, the model has the shortest convergence time and the fastest average convergence speed across all image resolutions, including 40X, 100X, 200X, and 400X. This further reinforces the superior performance and potential of the model for medical image classification tasks. These results proved the efficiency and effectiveness of the model.
@@ V Performance evaluation and visualization
To thoroughly oversee the model’s performance and glean insights into the training process, focus on several key metrics, including training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics can demonstrate the model’s training performance. The results of analyzing these key metrics are presented in detail below.
Refer to caption 40X
Refer to caption 100X
Refer to caption 200X
Refer to caption 400X
Fig. 7: Training and validation loss with error bars
Refer to caption 40X
Refer to caption 100X
Refer to caption 200X
Refer to caption 400X
@@ V-A Loss analysis
This study utilized a dual-activated lightweight attention ResNet50 model for breast cancer classification. The critical metrics for monitoring model performance and understanding the training process are training loss and validation loss [55, 56, 57]. These loss values play an important role when evaluating model performance, and their visualization gives us a clearer picture of the training progress.
Fig. 7 shows the variation of training loss and validation loss with the number of training iterations for different image resolutions (including 40X, 100X, 200X, and 400X). The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function. Line plots were used to show the trend of training loss and validation loss, and error bars for each data point were used to indicate the standard deviation of the loss values.
As training progresses, a gradual decrease in training loss and validation loss is observed, which signifies that the model is effectively learning and adapting to the data. This is a positive trend and shows that the model is making progress on the task. In the later stages of training, the validation loss changes less. It offers a stable trend, indicating that the model steers from overfitting on the validation set and demonstrates strong generalization to new data.
@@ V-B Accuracy analysis
In Fig. 8, the evolution of training accuracy and validation accuracy at different image resolutions (40X, 100X, 200X, and 400X) is depicted [58]. The horizontal axis represents the number of training iterations, while the vertical axis represents the accuracy. A line graph is employed to portray the accuracy trend, and error bars are incorporated to convey the standard deviation of the accuracy values.
@@ VI Summary and conclusions
This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets. The ResNet50 model incorporates the lightweight attention mechanism and a dropout layer to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of unbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more investigations of breast multiclassification problems are planned.

### v8

@@ III-B Dual-activated lightweight attention ResNet50
A new approach for automatic breast cancer classification is proposed in this research using a new version of the ResNet50 technique called Dual-activated lightweight attention ResNet50 (DALAResNet50). The DALAResNet50 model is methodically structured into three core stages: data preprocessing and augmentation, transfer learning based on ResNet50 architecture, and DALAResNet50 model optimization, as detailed below:
III-B1 Data preprocessing and augmentation
Initially, all breast cancer histopathology images are resized to 224x224 pixels, aligning with the input specifications for convolutional neural networks. A normalization process follows, adjusting pixel values to a range between [-1, 1] to expedite model training and augment the model’s generalization capability with novel data. Data augmentation strategies like random rotation and horizontal flipping bolster the model’s robustness. These strategies mimic potential real-world image alterations, improving the model’s resilience to subtle image variations.
III-B2 Transfer learning based on ResNet50
At this stage, capitalizes on a pre-trained ResNet50 framework, incorporating a custom-designed attention module within its fourth residual block. This module aims to pinpoint small-scale cell nuclei characteristic of breast cancer and glean pertinent semantic information. This strategic enhancement significantly boosts the model’s ability to learn and discern features.
III-B3 DALAResNet50 model optimization
During the optimization phase of the DALAResNet50 model, it is adopted that focuses sequentially on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability. The related steps are as follows:
  1. 1.
    Initially, to address the issue of spatial variations in input feature maps, traditional pooling layers are replaced with adaptive average pooling. This adjustment optimizes the network’s parameter structure, enabling the model to process global contextual information more effectively and accurately. The introduction of adaptive average pooling not only simplifies the network’s complexity but also enhances the model’s ability to capture important information from different parts of the image, laying a solid foundation for subsequent feature processing.
  2. 2.
    Next, the model employs a dual activation function strategy in the activation layer, starting with the LeakyReLU activation function [10] in the initial phase. This function is chosen for its unique capability to maintain a slight yet nonzero gradient when the unit is inactive, effectively combating the problem of vanishing gradients and significantly boosting the training efficiency of deep neural networks. Following this, the ReLU activation function [11] is applied in the subsequent fully connected layer to optimize model performance further and enhance the representation of image features. The careful combination of LeakyReLU and ReLU activation functions dramatically improves the model’s ability to process and learn complex image characteristics, providing strong support for accurately classifying breast cancer images.
  3. 3.
    Finally, the introduction of element-wise multiplication operations further enhances the model’s Feature Adaptability, allowing for dynamic, nonlinear adjustments of features based on specific inputs. This step significantly increases the network’s flexibility and information utilization, enabling the model to adapt its processing flow according to the particular circumstances of the input features. This approach notably improves the model’s feature processing capability, leading to higher accuracy and efficiency in classifying breast cancer images.
@@ III-B3 DALAResNet50 model optimization
Fig. 3 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively. These flowcharts highlight the newly added features in the DALAResNet50. A comprehensive comparative analysis of these differences is detailed in Table 2. The visualization of the DALAResNet50 approach is depicted in Fig. 4.
@@ IV-A BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathology images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed dataset statistics in Table 3 cover sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 5 presents these pathological images as a practical example.
40x
100x
200x
400x
@@
Benign 625 644 623 588
Malignant 1370 1437 1390 1232
Total No of Images 1995 2081 2013 1820
@@ IV-C Experiment implementation
In deep learning, the configuration of various parameters is paramount. Notably, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and imbalanced datasets to ensure that all categories are given proper attention. In addition, setting the training batch size and the number of iterations is significant. In addition, setting the training batch size and the number of iterations is substantial. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to balance training effectiveness and computational efficiency. Finally, this study also specifically considered the dropout rate setting. Previous studies generally set the dropout rate between 0.25 and 0.5. In processing the imbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability. Therefore, this study chose a lower dropout rate of 0.25.
The method proposed in this study and the benchmark model illustrated in Table 6 apply the above settings to ensure the experiments’ fairness and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
@@
SEResNet50 [39] 0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
DensNet121 [40] 0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
VGG16 [41] 0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
VGG16Inception [42] 0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
ViT [43] 0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
Swin-Transformer [44] 0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
Dinov2_Vitb14 [45] 0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
RseNet50 [9] 0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
@@
SEResNet50 [39] 0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
DensNet121 [40] 0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
VGG16 [41] 0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
VGG16Inception [42] 0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
ViT [43] 0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
Swin-Transformer [44] 0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
Dinov2_Vitb14 [45] 0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
RseNet50 [9] 0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
@@
SEResNet50 [39] 2696.22 2814.30 2711.89 2356.18 27073090 2644.65
DenseNet121 [40] 2352.46 2643.79 2423.36 2078.34 9028970 2347.49
VGG16 [41] 3212.95 3182.54 3011.53 2640.19 134268738 3011.80
VGG16Inception [42] 3074.99 3253.50 3303.12 2682.61 38139938 3078.56
ViT [43] 2932.11 3022.54 3013.86 2691.16 51375106 2914.92
Swin-Transformer [44] 3020.11 3132.42 2940.98 2592.96 86745274 2921.62
Dinov2_Vitb14 [45] 2895.02 2303.25 2245.98 2017.15 86582018 2365.35
ResNet50 [9] 2920.53 2933.26 2307.94 2097.33 21797672 2564.77
@@ IV-D Baseline selection
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze imbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. The model uses a pre-trained network to migrate learned weights from one domain to another to reduce training time and computational cost.
@@
40X
100X
200X
400X
Fig. 6: Training and validation loss with error bars
@@ IV-E Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], Imbalance Balanced Accuracy (IBA) [50], and Geometric Mean (GMean) [51], to assess the performance of the model holistically. Below is an introduction to these metrics:
  1. 1.
    Accuracy is one of the most intuitive and commonly used metrics for evaluating the overall performance of a model. It calculates the proportion of all samples (whether positive or negative) that were predicted correctly relative to the total number of samples [46, 47, 48].
  2. 2.
    The F1 Score is a crucial metric that combines precision and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly [52].
  3. 3.
    IBA is a metric for evaluating the performance of classification models on imbalanced datasets, adjusting standard accuracy to more fairly account for the model’s ability to identify both majority and minority classes, ensuring the performance assessment reflects classification effectiveness across all categories. [53].
  4. 4.
    GMean is a metric that calculates the geometric mean of individual class accuracies, indicating a model’s capability to address category imbalance. It effectively equalizes the performance assessment for both dominant and infrequent categories, emphasizing the model’s comprehensive efficiency in handling variances in class distribution. This metric is in scenarios where class imbalance could significantly skew the performance metrics, ensuring that the model’s ability to classify less represented classes correctly is adequately considered in its overall [54].
Table 6 showcases the model’s performance on the test set at various image resolutions (40X, 100X, 200X, and 400X). Across these resolutions, the model consistently outperforms nine other popular networks regarding accuracy, F1 score, IBA, and GMean. This consistency underlines the network’s significant advantage and underscores its potential for medical image classification tasks.
Table 7 compares the model against eight others regarding parameter count and convergence speed. While DenseNet-121 has the fewest parameters, the proposed model achieves the shortest convergence times, exhibiting the fastest average speed across all resolutions (40X, 100X, 200X, and 400X). This further reinforces the proposed model’s efficiency and strong potential for medical image classification, highlighting superior performance and effectiveness.
@@ V Performance evaluation and visualization
To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics offers a comprehensive view of the model’s training behavior. An in-depth analysis of these key metrics and their implications is outlined below:
40X
100X
200X
400X
@@ V-A Loss analysis
This study deployed a DALAResNet50 model for the classification of breast cancer. Critical metrics for assessing model performance and understanding the training dynamics include training loss and validation loss [55, 56, 57]. These metrics are crucial for evaluating the model’s effectiveness, and their graphical representation offers a detailed insight into the training progression.
Fig. 6 illustrates the fluctuations in training and validation losses across varying numbers of training iterations at different image magnifications (40X, 100X, 200X, and 400X). The x-axis represents the number of training iterations, while the y-axis reflects the loss values. Line plots demonstrate the trends in training and validation losses, with error bars on each point indicating the standard deviation of these values.
Training and validation losses consistently decline, indicating the model’s successful learning and adaptation to the dataset. This downward trend is encouraging, as it signals ongoing improvement in task performance. In the later training phases, the validation loss stabilizes, suggesting that the model avoids overfitting on the validation dataset and exhibits robust generalization capabilities to unseen data.
@@ V-B Accuracy analysis
In Fig.7, the progression of training and validation accuracy across various image resolutions (40X, 100X, 200X, and 400X) is illustrated [58]. The horizontal axis indicates the number of training iterations, whereas the vertical axis displays accuracy levels. A line graph is utilized to depict the trend in accuracy, with error bars included to illustrate the standard deviation of the accuracy figures.
@@ VI Summary and conclusions
This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets. The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more investigations of breast multiclassification problems are planned.